OpenStudy (anonymous):

Given a = -1, h = 4, and k = -2, what is the equation of the graph if the parent function is y equals the square root of x?

OpenStudy (anonymous):

\[y = \sqrt{x}\]

OpenStudy (anonymous):

\[y = a \sqrt{x- h} + k\]

OpenStudy (anonymous):

\[y = -\sqrt{x-4}-2\]